OpenStudy (anonymous):

derive ((cos x)/(1-sinx))^2 thanks

OpenStudy (agent0smith):

\[\Large \frac{ \cos^2x }{(1-\sin x ) ^2 } =\] what are we supposed to do? Simplify? Differentiate?